### v2

# Simulating fermionic fractional Chern insulators with infinite projected entangled-pair states

Abstract
Infinite projected entangled-pair states (iPEPS) provide a powerful variational framework for two-dimensional quantum matter and have been widely used to capture bosonic topological order, including chiral spin liquids. Here we extend this approach to fermionic topological order by variationally optimizing $U(1)$-symmetric fermionic iPEPS for a fractional Chern insulator (FCI), with bond dimensions up to $D=9$. We find evidence for a critical bond dimension, above which the ansatz faithfully represents the FCI phase. The FCI state is characterized using bulk observables, including the equal-time single-particle Green’s function and the pair-correlation function, as well as the momentum-resolved edge entanglement spectrum. To enable entanglement-spectrum calculations for large iPEPS unit cells, we introduce a compression scheme and show that the low-lying part of the spectrum is already well converged at relatively small cutoff dimensions.
preprint: APS/123-QED
Introduction—Fractional Chern insulators (FCIs) are lattice analogues of the fractional quantum Hall (FQH) states realized without external magnetic fields [33, 47, 48] in fractionally filled narrow Chern bands [51, 52]. As a chiral topologically ordered phase, FCIs host fractionalized anyonic excitations and chiral gapless edge modes. Recent experimental realizations of this phase of matter in two-dimensional (2D) moiré materials [5, 40, 57, 29] have reignited the interest in FCIs. Simulating microscopic Hamiltonians that host FCIs is therefore crucial—both to test theory and to connect with experimental observations. So far, the primary numerical methods employed in this field have been restricted to exact diagonalization (ED) [33, 47, 48, 2] and infinite density matrix renormalization group (iDMRG) [22, 25], which are limited to small clusters or finite-width cylinders and hence are subjected to finite-size effects.
Infinite projected entangled-pair state (iPEPS), a two-dimensional tensor network ansatz, offers a complementary route: by construction, it is formulated directly in the thermodynamic limit. Despite the no-go theorem forbidding an exact finite-correlation-length iPEPS representation of a gapped chiral phase [15], iPEPS has proven effective in practice for simulating bosonic chiral topological orders, including chiral spin liquids (CSLs) [41, 7, 24, 36, 4, 8], bosonic FQH states [55], and non-Abelian generalizations [9, 37]. Nevertheless, iPEPS has not yet been applied to fermionic topological order to the best of our knowledge.
Recent algorithmic advances have made iPEPS a powerful toolkit for tackling this problem. First, we leverage standard $U(1)$-symmetric iPEPS [30, 49, 1, 50] to implement fermionic charge conservation, which enables efficient encoding of fermionic statistics [14, 11] and also yields block-sparse tensors that substantially reduce computational cost, enabling optimizations at larger bond dimensions. Second, we use automatic differentiation (AD) [28] that facilitates straightforward gradient-based optimizations. Unlike the standard imaginary–time evolutions, the gradient-based approach can reliably find long-range entangled phases such as CSLs and the FCI studied here. The key new advance comes from the fixed-point AD method [28, 42, 32, 17]. When combined with a robust gauge-fixing scheme [17], we can overcome the performance bottleneck in simulating chiral states: the corner transfer matrix renormalization group (CTMRG) [34, 35, 39, 13, 16], often adopted to compute observables for iPEPS, regularly requires a large number of iterations ($O(10^{3})$ for the FCI state studied here) to converge due to the long correlation lengths implied by the no-go theorem. This leads to large memory requirements when backpropagating by AD. Fixed-point AD circumvents this by directly differentiating the converged fixed point of CTMRG and eliminates the need for storing any intermediates throughout CTMRG.
In this Letter, we extend the applicability of iPEPS to fermionic FCIs by integrating the above advances. We carry out gradient optimization on a $U(1)$-symmetric fermionic iPEPS for a spinless-fermion Hamiltonian on the honeycomb lattice and are able to the push the bond dimension of iPEPS to $D=9$. We find that a faithful representation of the FCI requires $D\geq 7$, and identify the phase via the equal-time single-particle Green’s function, the pair-correlation function, and the edge entanglement spectrum (ES). To access the ES for large iPEPS unit cells, we further introduce a simple compression scheme, which makes the calculations computationally feasible.
[…]
Model and methods—We study spinless fermions on the honeycomb lattice governed by the Haldane model Hamiltonian [23] with density-density interactions
$\hat{H}=-\sum_{i,j}t_{ij}\hat{c}_{i}^{\dagger}\hat{c}_{j}+V_{1}\sum_{\langle ij\rangle}\hat{n}_{i}\hat{n}_{j},$ (1)
where $\langle ij\rangle$ denotes nearest-neighbor (NN) sites, $\hat{c}^{\dagger}_{i}$ creates a spinless fermion on site $i$ of the honeycomb lattice and $\hat{n}_{i}=\hat{c}^{\dagger}_{i}\hat{c}_{i}$ measures the site occupation number. Fermions can hop between NN, second-NN, and third-NN sites with hopping amplitudes $t_{1},t_{2}e^{\pm i\phi},$ and $t_{3}$, respectively. The sign convention of the acquired phase $\pm\phi$ is illustrated in Fig. 1(a). We focus on $V_{1}/t_{1}=10.0,t_{2}/t_{1}=0.7,t_{3}/t_{1}=-0.9$, and $\phi=0.35\pi$, for which the non-interacting part of the Hamiltonian features an almost flat Chern band as the valence band. When this band is fractionally filled to $\nu=1/3$ filling (one fermion per three unit cells; here $\nu=2$ denotes full lattice filling, i.e., one fermion per honeycomb site) the NN interactions in the Hamiltonian give rise to an FCI ground state, as evidenced by iDMRG simulations [22] on thin cylinders. The FCI phase persists upon including second-NN and third-NN interactions [6].
Here, we simulate the ground state of this model directly in the thermodynamic limit, by variationally optimizing an iPEPS. The iPEPS encodes the amplitudes $\phi_{\vec{\nu}}$ of the ground state $|\psi\rangle=\sum_{\vec{\nu}}\phi_{\vec{\nu}}\prod_{i}(\hat{c}_{i}^{\dagger})^{\nu_{i}}|0\rangle$ with $\nu_{i}\in\{0,1\}$ as a contraction of an infinite, translationally invariant two-dimensional tensor network. To construct our iPEPS, we first coarse-grain the honeycomb lattice into its Bravais lattice as shown in Fig. 1(a), by assigning a local physical Hilbert space $\mathbb{V}_{\text{phy}}=\text{span}(|0\rangle,\hat{c}^{\dagger}_{\mathrm{A}}|0\rangle,\hat{c}^{\dagger}_{\mathrm{B}}|0\rangle,\hat{c}^{\dagger}_{\mathrm{A}}\hat{c}^{\dagger}_{\mathrm{B}}|0\rangle)$ of dimension $d=4$ to each unit cell. This is equivalent to a two-flavour fermion with one flavour for each sublattice. Next, we assign a rank-5 tensor $(A^{[\bm{r}]})_{i_{t}i_{l}i_{b}i_{r}i_{p}}$ to each Bravais lattice unit cell $\bm{r}$
$A^{[\bm{r}]}:\mathbb{V}_{t}\otimes\mathbb{V}_{l}\otimes\mathbb{V}_{b}^{*}\otimes\mathbb{V}_{r}^{*}\otimes\mathbb{V}{{}_{\text{phy}}}\to\mathbb{C},$ (2)
where $\mathbb{V}_{\alpha=t,l,b,r}$ are virtual Hilbert spaces of bond dimension $D$ associated with the corresponding legs of $A^{[\bm{r}]}$, and $\mathbb{V}^{*}$ denotes the dual space of $\mathbb{V}$. Finally, the tensors $A^{[\bm{r}]}$ are contracted through their virtual legs $(t,l,b,r)$ according to the lattice connectivity, forming the iPEPS shown in Fig. 1(b). This ansatz efficiently approximates amplitudes $\phi_{\vec{\nu}}$, with the precision controlled by $D$.
[…]
Observables, in particular the single-particle Green’s function $\langle\hat{c}^{\dagger}_{i}\hat{c}_{j}\rangle$ and density correlations $\langle\hat{n}_{i}\hat{n}_{j}\rangle$, are evaluated via the contractions of the corresponding infinite double-layer networks. Here, we use the directional CTMRG algorithm [13, 16], which iteratively finds a set of finite corner and edge environment tensors $\{C,T\}$ that approximates the semi-infinite surroundings of any finite rectangular patch of the double layer network supporting the observables [Fig. 6(a,b)]. Contracting these environments with the patch yields the desired expectation values, with an accuracy controlled by the environment bond dimension $\chi$ of $\{C,T\}$; the exact results are recovered in the $\chi\to\infty$ limit. Finally, to find the optimal iPEPS describing the ground state, we minimize the energy density $e_{0}$ via gradient-based L-BFGS optimization of the tensor elements in $\vec{A}$, combined with a strong Wolfe line search [38]. Crucially, the gradient $\mathrm{d}e_{0}/\mathrm{d}\vec{A}$ is computed via the fixed-point method [28, 43, 17], eliminating memory bottleneck in plain backpropagation due to $O(10^{3})$ CTMRG steps required for convergence. For further technical details of the tensor network methods used in this work, we refer to Appendices A and B.
Figure 2: (a) Ground-state energy density $e_{0}(D,\chi)$ of the optimized iPEPS for bond dimensions from $D=4$ to $D=9$ and its extrapolations to $\chi\to\infty$. The extrapolated values are plotted in the inset as a function of $D$. (b) Variance of the real-space charge distribution of the optimized states. For comparison, we also show the corresponding charge profiles $\{\langle\hat{n}_{\alpha,\bm{r}}\rangle\}$ for $D=4$ and $D=9$, plotted with a common color scale.
Results—We optimize the $U(1)$-symmetric iPEPS up to $D=9$, choosing an identical $U(1)$ structure on all virtual spaces $\mathbb{V}_{t,l,b,r}=\oplus_{q}\mathbb{V}_{q}$, which we summarize in Table 1. First, we focus on the energetics of the optimized iPEPS, shown in Fig. 2(a), together with the extrapolations of the ground state energy density to the $\chi\to\infty$ limit, which obey variational principle, serving as strict upper bounds on the true ground-state energy density in the thermodynamic limit. As shown in the inset, the energy decreases substantially when increasing bond dimension from $D=4$ to $D=7$. Further increasing $D$ yields only a modest improvement. Moreover, by comparing with the ED results on finite torus clusters (up to 24 unit cells), we find that the iPEPS energy for $D\leq 6$ lies above the first excitation gap, indicating strong competition among states within the variational manifolds of $D\leq 6$, while $e_{0}$ of $D\geq 7$ are all well below the first excited gap. A similar behavior is found for real-space charge inhomogeneity. We quantify it by the variance of the local density within the iPEPS unit cell [Fig. 2(b)],
$\text{Var}(\langle\hat{n}_{\alpha,\bm{r}}\rangle)=\frac{\sum_{\bm{r},\alpha=\text{A,B}}(\langle\hat{n}_{\alpha,\bm{r}}\rangle-\bar{n})^{2}}{N},$ (3)
where the sum runs over honeycomb lattice sites in the iPEPS unit cell, $N$ is the number of sites in the iPEPS unit cell ($N=18$ in our case), and $\bar{n}=1/6$ (per site) for $\nu=1/3$. We find that the $D<7$ states are also associated with significantly larger charge inhomogeneity. These results imply the existence of a minimal bond dimension required to accurately capture the ground state and its chiral topological order, as also observed in bosonic cases [9, 36]. For $D<D_{\text{min}}=7$, the variational manifold favors states with charge modulations and without clear entanglement-spectrum signatures of an FCI state. In the remainder of the paper, we therefore focus on the states with $D\geq 7$.
Figure 3: Equal-time single-particle Green’s function $|\langle\hat{c}_{\mathrm{A},0}\hat{c}^{\dagger}_{\mathrm{A},\bm{r}}\rangle|$ between A-A sublattice sites. Distance $|\bm{r}|$ is calculated in Cartesian coordinates and $a$ is the honeycomb lattice constant. (a) Correlations along the two primitive lattice vector directions $\bm{a_{1}}$ and $\bm{a_{2}}$, evaluated for the $D=9$ state with $\chi=468$. (b) The Green’s function along $\bm{a_{2}}$ of the $D=9$ state with increasing $\chi$.
Next, we discuss the bulk properties captured by the correlation functions. In Fig. 3(a), we present the equal-time single-particle Green’s function $\langle\hat{c}_{\text{A,0}}\hat{c}^{\dagger}_{\text{A},\bm{r}}\rangle$, evaluated along the two primitive lattice vectors $\bm{{a}_{1}}$ and $\bm{a}_{2}$. We observe a similar behavior as found in CSLs [41, 24, 36, 4, 8] and bosonic FQH states [55]: a rapid exponential decay with a correlation length $\xi_{\mathrm{bulk}}/a=0.63$ ($a$ is the honeycomb lattice constant)—consistent with a gapped bulk—followed by an artifact of iPEPS representation, a long-range “gossamer” tail (note the relatively tiny amplitude of the tail). The latter reflects the no-go theorem, prohibiting exact finite-$D$ iPEPS representation of states with chiral edge modes [15]. Increasing $\chi$ flattens the long-range tail but leaves the bulk part unchanged [see Fig. 3(b)]. In practice, the no-go theorem does not obstruct the identification of the key FCI signatures.
The pair-correlation function $g(\bm{x})$ is another important quantity in FQH physics [21, 20]. It reveals information about the probability of finding two fermions at a relative distance $\bm{x}$. The exponentially converging behavior of $g(\bm{x})$ to unity as $|\bm{x}|/\ell_{B}\to\infty$ is a signature of incompressibility ($\ell_{B}$ is the magnetic length). At short distances $|\bm{x}|/\ell_{B}\to 0$, $g(\bm{x})$ exhibits a correlation hole, reflecting the generalized exclusion principle of FQH states. On the lattice, $g(\bm{x})$ is defined as
$g(\bm{x})=\frac{\langle\psi|\hat{n}_{0}\hat{n}_{\bm{x}}|\psi\rangle}{\langle\psi|\hat{n}_{0}|\psi\rangle\langle\psi|\hat{n}_{\bm{x}}|\psi\rangle},$ (4)
where $\hat{n}_{\bm{x}}$ is the local fermion number operator at site $\bm{x}$. On the honeycomb lattice, $\hat{n}_{\bm{x}}$ acts on either A or B sublattice depending on the position $\bm{x}$. In Fig. 1(d), we show the pair-correlation function of the FCI state with the reference site fixed to the B sublattice at the origin. Correlations along the two primitive lattice vectors $\bm{a_{1}}$ and $\bm{a_{2}}$ are shown and nearly coincide. The function $g(\bm{x})$ exhibits short-range oscillations and rapidly approaches unity. For comparison, Fig. 1(c) shows $g(\bm{x})$ for the corresponding CI obtained by simulating the model with $V_{1}=0$ at $\nu=1$. In this case $g(\bm{x})$ approaches unity monotonically without oscillations. Furthermore, we overlay continuum $g(\bm{x})$ for the Laughlin $\nu=1/3$ state (via the polynomial-expansion method [18]) and for the non-interacting $\nu=1$ IQH state, $g(\bm{x})=1-\exp(-|\bm{x}|^{2}/2\ell_{B}^{2})$ in the lowest Landau level. Remarkably, we find very good agreement with the lattice data for both cases using a single magnetic length $\ell_{B}/a\approx 1.63$. This result underscores how even quantitative properties of these states transcend the particular microscopic realization.
Table 1: Optimized $U(1)$-symmetric iPEPS. Total bond dimension $D=\sum_{q}D_{q}$ and bond dimensions $D_{q}=\text{dim}(\mathbb{V}_{q})$ of the individual charge-$q$ sectors, largest environment bond dimension $\chi_{\text{opt}}$ used during optimizations, and variational estimate of ground-state energy density $e_{0}(\chi\to\infty)$ obtained by extrapolations in $\chi$ [see Fig. 2(a)].
$D$ $\{D_{q}\}$ $\chi_{\text{opt}}$ $e_{0}(\chi\to\infty)$
4 $(D_{-1},D_{0},D_{1})=(1,2,1)$ 64 -0.8444092(17)
5 $(D_{-1},D_{0},D_{1})=(1,2,2)$ 75 -0.8514863(5)
6 $(D_{-1},D_{0},D_{1})=(2,2,2)$ 72 -0.855254(5)
7 $(D_{-1},D_{0},D_{1})=(2,3,2)$ 98 -0.8592989(18)
8 $(D_{-1},D_{0},D_{1})=(2,3,3)$ 112 -0.859490(4)
9 $(D_{-1},D_{0},D_{1})=(3,3,3)$ 117 -0.859614(12)
Turning to truly universal properties, chiral topological order of FCIs can be characterized by their gapless edge theory, which, in our case, is a chiral $U(1)$ boson theory. Importantly, due to the Li–Haldane correspondence [27, 45], we do not need to impose open boundaries to create a physical edge, instead, the low-lying edge spectrum is obtained from the entanglement spectrum (ES) of the ground state, which is readily accessible within iPEPS via the bulk–edge correspondence [10]. More specifically, we use the optimized iPEPS $|\psi\rangle$ to construct a PEPS wave function $|\Psi\rangle$ on an infinite horizontal cylinder of width $W$ (in unit cells) by tiling the cylinder with iPEPS unit cells. Under the bipartition of the cylinder into left and right halves, the entanglement Hamiltonian $H_{\text{ent}}$ is defined as the logarithm of the reduced density matrix for the left part $H_{\text{ent}}=-\ln(\rho_{\text{L}})$ with $\rho_{\text{L}}=\text{Tr}_{\text{R}}|\Psi\rangle\langle\Psi|$ (or analogously for the right half). As shown in Ref. [10], the eigenvalues of $\rho_{\text{L}}$ can be calculated using the following relation
$\text{spec}(\rho_{\text{L}})=\text{spec}(\sqrt{\sigma_{\text{L}}^{T}}\sigma_{\text{R}}\sqrt{\sigma_{\text{L}}^{T}})=\text{spec}(\sigma_{\text{L}}^{T}\sigma_{\text{R}}),$ (5)
where $\text{spec}(A)$ denotes the non-zero spectrum of the matrix $A$, and $\sigma_{\text{L}}$ and $\sigma_{\text{R}}$ are matrices formed from the left and right leading eigenvectors of the column-to-column transfer matrix of the double-layer norm $\langle\Psi|\Psi\rangle$ network. Since we use a $3\times 3$ iPEPS unit cell, the transfer matrix consists of three columns. The matrices $\sigma_{\mathrm{L}}$, $\sigma_{\mathrm{R}}$, and their product $\sigma_{\mathrm{L}}^{T}\sigma_{\mathrm{R}}$ are approximated by the matrix–product operators (MPOs) constructed from the CTMRG environment tensors, as illustrated in Fig. 4(a,b). Note that swap gates are introduced to encode fermionic statistics, which can be subsequently absorbed into the MPO tensors. For a fermionic PEPS on a finite-width cylinder one may impose anti-periodic (APBC) or periodic (PBC) boundary conditions along the circumference, depending on the definition of the lattice translation symmetry (see Appendix C). In Fig. 4(b) we have implicitly assumed APBC.
Figure 4: (a) Left/right dominant eigenvectors $\sigma_{\mathrm{L}}$ and $\sigma_{\mathrm{R}}$ of the column-to-column double-layer transfer matrix. For clarity, we omit the arrow directions on tensor legs. The eigenvectors are approximated by MPOs assembled from the CTMRG environment edge tensors $T_{\mathrm{L}/\mathrm{R}}^{[i]}$; see Fig. 6 for the definitions of the double-layer tensor and the environment tensors. (b) MPO representation of $\sigma_{\mathrm{L}}^{T}\sigma_{\mathrm{R}}$ on a cylinder of width $W=2$. The MPO for $\sigma_{\mathrm{L}}^{T}\sigma_{\mathrm{R}}$ has a three-tensor unit cell; small squares indicate fermionic swap gates. (c) MPO unit-cell compression for a 3-site unit cell: at each step, isometries $P_{\mathrm{L}}^{(i)}$ and $P_{\mathrm{R}}^{(i)}$ block two neighboring sites within the unit cell, reducing the unit-cell size by one while truncating the MPO “physical” leg to dimension $\leq D_{\mathrm{c}}$.
Directly solving for the spectrum of the periodic MPO $\sigma_{\text{L}}^{T}\sigma_{\text{R}}$ quickly becomes infeasible as the cylinder width $W$ increases, especially for large MPO unit cells: The Hilbert-space dimension grows as $O(D^{WM})$, where $D$ is the “physical” dimension of the MPO (equal to the iPEPS bond dimension), and $M$ is the MPO unit-cell size. We therefore introduce a pre-compression step before diagonalization. The idea is to coarse-grain the MPO unit cell to a single site, reducing the dimension from $D^{M}$ to an effective dimension $D_{\text{c}}$. The procedure is summarized in Fig. 4(c). At step-$i$, we apply isometries $P_{\text{L/R}}^{(i)}$ to two neighboring sites within the unit cell and obtain a single site with a dimension bounded by $D_{\text{c}}$. The iteration is repeated until the unit cell is coarse grained into a single site. The construction of $P^{(i)}_{\text{L/R}}$ follows Ref. [56], which utilizes the higher-order singular value decomposition (HOSVD). For simplicity, we always construct the isometries based on the HOSVD of $\sigma_{\text{L}}$ and then apply the isometries to both $\sigma_{L}$ and $\sigma_{R}$.
Figure 5: Momentum-resolved ES of the $D=7$ state on a cylinder of circumference $W=5$ (15 sites) under APBC. The spectrum is calculated for the $U(1)$ charge sector $n=0$. (a) The cutoff bond dimension $D_{\mathrm{c}}=35$ is fixed, and the environment bond dimension $\chi$ is varied. (b) The environment bond dimension $\chi=35$ is fixed, and the cutoff bond dimension $D_{\mathrm{c}}$ is varied. In both cases, the low-level counting across momentum sectors follows $(1,1,2,3,5)$, consistent with the chiral $U(1)$ boson edge of the $\nu=1/3$ Laughlin state.
With the pre-compression step, we are able to push the cylinder width to $W=5$ (15 sites) for the $D=7$ iPEPS. Figure 5 shows the resulting momentum-resolved ES in the $U(1)$ charge sector $n=0$ under APBC and compares the spectra obtained at several environment bond dimensions $\chi$ and MPO cutoff dimensions $D_{\text{c}}$. Remarkably, despite coarse-graining an MPO unit cell of dimension $D^{M}=7^{3}=343$ to a single site of dimension $D_{\text{c}}\approx 28\text{–}42$, the low-energy levels are robust across these parameters and are already well converged even for a modest $\chi$ and $D_{\text{c}}$ (e.g., $\chi=28$ and $D_{\text{c}}=35$). Most importantly, we observe the characteristic level counting $(1,1,2,3,5)$ for $W=5$, providing a sharp fingerprint of the FCI state [27, 45]. The same chiral $U(1)$ boson level counting is obtained under PBC, and it persists for the $D=8$ and $D=9$ states, as shown in the Appendix C.
Discussion—By combining state-of-the-art techniques, we have demonstrated that iPEPS can simulate chiral topological order in fermionic systems by variationally realizing an FCI at filling $\nu=1/3$. In the bulk, the equal-time single-particle Green’s function shows a rapid exponential decay followed by a long-range “gossamer” tail—consistent with the finite-$D$ iPEPS manifestation of the no-go constraint—and the pair-correlation function $g(\bm{x})$ matches qualitative features of the $\nu=1/3$ Laughlin state, including the correlation hole and the rapid convergence to unity. The edge ES exhibits the characteristic $(1,1,2,3,5)$ counting. It is already well-converged at modest environment bond dimension $\chi$ and compression cutoff $D_{\text{c}}$, and remains robust to the choice of APBC/PBC along the cylinder.
Our methodology readily extends to other filling fractions and lattice geometries, which would be useful in studying states in the hierarchy of composite fermion states and non-Abelian FCIs. The remaining computational bottleneck in simulating chiral topological order is the slow convergence of CTMRG in the forward pass (the backward pass in the fixed-point approach typically converges much faster). A fast algorithm for determining the environment tensors is highly desired. In this work we fix the fractional filling by using a $U(1)$-symmetric ansatz and a multi-site iPEPS unit cell. Since FCI states do not break translation symmetry, a natural question is whether a single-site iPEPS could approximate a fermionic FCI by relaxing $U(1)$ to $\mathbb{Z}_{2}$ and tuning the density via a chemical potential. Single-site iPEPS are easier to optimize and can be further endowed with lattice symmetries. In addition to the ground state, another interesting direction is to apply recently developed iPEPS excitation ansatz [54, 44, 42, 53] to compute neutral and charged excitations on top of the FCI state, providing dispersion relations and gaps directly in the thermodynamic limit.
Our work adds iPEPS to the few numerical methods that can be used to simulate chiral fermionic topological order, in a time where experimental realizations of such exotic phases call for quantitatively reliable predictions.
Acknowledgments—We thank Marek Rams, Anna Fracuz, Paul Brehmer, Johannes Motruk, and Sen Niu for valuable discussions. Tensor network simulations were performed with open-source library YASTN [46].
The authors acknowledge support from the Swiss National Science Foundation through a Consolidator Grant (iTQC, TMCG-2213805) and a Quantum grant (20QU-1225225). This project further received financial support from Google. We thank NVIDIA for providing access to GPU computing resources that supported this work.